ASAP
ostct Financlals Using Ratios Professor Michael Kinnen FIN331 Question2 Otto is the CEO of Heinz Sports Partners (HSP), and the firm has 2022 revenue of $10.6 million. Its net income margin is 25%. Use the information below to a) What is the firms net income? b) If the firms ROE is 27%, what is the total amount of equity for this company c) If the firm has a debt-equity ratio of 70%, how much debt does HSP have? d) What is the total assets for this firm? e) HSPs current liabilities are 45% of the total liabilities, what is the balance of current liabilities and 0) g) also for long term liabilities? The firm has net working capital of $3,850,000. Calculate HSPs current asset balance. How much does the firm have in fixed assets? h) HSPs common stock has a book value of $1.35 per share. If there are 2,140,000 shares outstanding how much is the total balance for common stock? What is the total retained earnings?
Answer #1

Answer:

1)

Net income= margin * revenue=25%*10.6 million=$2.65 millions

2) ROE=27%

ROE= net income/Equity

Equity = Net Income / ROE=2.65 millions/ 27%=$9.8148 millions

3)

Debt-Equity Ratio=70%

Debt-equity ratio=Debt/equity

Debt=Equity* debt equity ratio=70%*9.8148=$6.87 millions

4)

Total Asset= Equity + Debt=9.8148+6.87=$16.6848 millions

5)

Current Liability = 45%*total liability=45%*6.87=$3.0915 millions

Long term liability = (1-45%)*total liability =55%*6.87=$3.7785 millions

6)

Current asset = Current Liability + Working Capital= 3.0915+3.85=$6.9415 millions

7)

Fixed asset= Total asset - Current Asset=16.6848-6.9415=$9.7433 millions

8)

Total value of common stock= share price * shares outstanding =1.35*2.14 =$2.889 millions

Total Retained earnings=Total Equity - Total value of common stock=9.8148-2.889=$6.9258 millions